if the lenght is 2 more than 4 times the width. the perimeter is 158.
is there something missing, i took it as saying the length is 4 times the width plus 2. is there a formula to break down perimeter backwards?
i came up with 15.4 as width, and 63.6 as lenght,to give me 158 for perimeter, but mainly thru trial and error method, i took a number times 4 then added 2 till i got 158, i know that wasnt the way it supposed to be done, was wondering what the formula and how work should be shown.

Answer by Alan3354(69443)

if the lenght is 2 more than 4 times the width. the perimeter is 158.
is there something missing, i took it as saying the length is 4 times the width plus 2. is there a formula to break down perimeter backwards?
i came up with 15.4 as width, and 63.6 as lenght,to give me 158 for perimeter, but mainly thru trial and error method, i took a number times 4 then added 2 till i got 158, i know that wasnt the way it supposed to be done, was wondering what the formula and how work should be shown.
------------------
L = 4W + 2
Perimeter = 2L + 2W = 158
2(4W+2) + 2W = 158
4W+2 + W = 79
5W = 77
W = 15.4
---------
L = 63.6
----------
P = 2L + 2W is the only "formula" involved
PS It's length, not lenght